Multiple Production Department Factory Overhead Rates Spotted Cow Dairy Company manufactures three products—whole milk, skim milk,...

  1. Multiple Production Department Factory Overhead Rates

    Spotted Cow Dairy Company manufactures three products—whole milk, skim milk, and cream—in two production departments, Blending and Packing. The factory overhead for Spotted Cow Dairy is $380,300.

    The three products consume both machine hours and direct labor hours in the two production departments as follows:

    Direct Labor Hours Machine Hours
    Blending Department
    Whole milk 610 830
    Skim milk 390 670
    Cream 300 530
    1,300 2,030
    Packing Department
    Whole milk 180 1,190
    Skim milk 190 730
    Cream 175 780
    545 2,700
    Total 1,845 4,730

    The management of Spotted Cow Dairy Company now plans to use the A method that allocates factory overhead to product by using factory overhead rates for each production department.multiple production department factory overhead rate method. The total factory overhead associated with each department is as follows:

    Blending Department $221,000
    Packing Department 159,300
    Total $380,300

    Required:

    If required, round all amounts to the nearest dollar.

    1. Determine the multiple production department factory overhead rates, using direct labor hours for the Blending Department and machine hours for the Packing Department.

    Blending Department $ per direct labor hour
    Packing Department $ per machine hour

    2. Determine the product factory overhead costs, using the multiple production department rates in (1).

    Whole milk Skim milk Cream
    Blending Department factory overhead $ $ $
    Packing Department factory overhead
    Total factory overhead $ $ $
Answer #1
1
Blending Department 170 per direct labor hour =221000/1300
Packing Department 59 per machine hour =159300/2700
2
Whole milk Skim milk Cream
Blending Department factory overhead 103700 66300 51000
Packing Department factory overhead 70210 43070 46020
Total factory overhead 173910 109370 97020
Workings:
Whole milk Skim milk Cream
Blending Department factory overhead =610*170 =390*170 =300*170
Packing Department factory overhead =1190*59 =730*59 =780*59